# Phenomenology of laminar acoustic streaming jets

### v1

@@ A.4 Normalization and dimensionless forms of the equations for the acoustic forcing field
$\begin{split}\widetilde{I}_{x}(x,r=0)=&\frac{1}{4}e^{-\frac{N}{L}\left(\sqrt{x^{2}+\frac{1}{4}}+x\right)}\Bigg{(}\left[\frac{x}{\sqrt{x^{2}+\frac{1}{4}}}+1\right]\left(\cosh\left[\frac{N}{L}\left(\sqrt{x^{2}+\frac{1}{4}}-x\right)\right]-\cos\left[\frac{2.44\pi}{S}\left(\sqrt{x^{2}+\frac{1}{4}}-x\right)\right]\right)\\ &+\left[1-\frac{x}{\sqrt{x^{2}+\frac{1}{4}}}\right]\sinh\left[\frac{N}{L}\left(\sqrt{x^{2}+\frac{1}{4}}-x\right)\right]+\frac{NS}{2.44\times\pi L}\left[1-\frac{x}{\sqrt{x^{2}+\frac{1}{4}}}\right]\\ &\times\sin\left[\frac{2.44\pi}{S}\left(\sqrt{x^{2}+\frac{1}{4}}-x\right)\right]\Bigg{)}\,,\end{split}$ (35)
@@ Appendix B Grid sensitivity analysis
Table 2 shows the effect of $p$-refinement on the position $x_{p}$ and amplitude $u_{x,p}$ of the velocity peak on the jet axis, for the case of the highest forcing magnitude investigated. The errors on both $x_{p}$ and $u_{x,p}$ fall below $0.1$ for $N_{p}\geq 6$. Despite their non-monotonic evolution with $N_{p}$, caused by $\widetilde{\bm{I}}$ being undersampled in the acoustic near field, these errors remain extremely small for further-refined grids. The results shown in the reminder of this document were obtained for $N_{p}=8$, for which the effects of the acoustic near field undersampling on the flow are proven to be negligible.

### v2

@@ Appendix B Grid sensitivity analysis
Table 2 shows the effect of $p$-refinement on the position $x_{p}$ and amplitude $u_{x,p}$ of the velocity peak on the jet axis, for the case of the highest forcing magnitude investigated. The errors on both $x_{p}$ and $u_{x,p}$ fall below $0.1$ % for $N_{p}\geq 6$. Despite their non-monotonic evolution with $N_{p}$, caused by $\widetilde{\bm{I}}$ being undersampled in the acoustic near field, these errors remain extremely small for further-refined grids. The results shown in the reminder of this document were obtained for $N_{p}=8$, for which the effects of the acoustic near field undersampling on the flow are proven to be negligible.